OpenStudy (anonymous):

Find the indicated missing measurement. Include the second solution if applicable. What am I doing wrong?

OpenStudy (anonymous):



OpenStudy (anonymous):

So this is what I have so far: sinA/a = sinB/b sin34 degrees/9 = sinB/7 sinB = 7sin34 degrees/9 sinB = 25.8 degrees Angle A is less than 90 degrees -> a is greater than or equal to bsinA bsinA = 7sin34 = 3.91 a is greater than or equal to bsinA = 2 solutions B' = 180-25.8 = 154.2 degrees But when I find for angle C of the second triangle it doesn't add up to 180... What am I doing wrong?